# Synthetic Geology - Structural Geology Meets Deep Learning

### v1

@@ Discussion
Our results demonstrate the feasibility and potential of using generative artificial intelligence (AI) models, particularly flow matching techniques, to construct realistic and geologically plausible three-dimensional structural models. The work presented here integrates two key components. First, we construct a process for creating a large, synthetic, and geologically consistent dataset that seeks to embed the rich and complex knowledge of structural geology into a mathematical framework suitable for training machine learning models. Second, we train a generative AI model that implicitly learns both unconditional and conditional probability distributions of geological structures. By doing so, it efficiently captures essential characteristics of the Earth’s crust and provides a highly complex prior to tackle the challenges common in geoscience applications, including data sparsity, uncertainty quantification, and the integration of diverse data modalities.
Advancements in Generative Modeling for Structural Geology.
@@
Generative AI for structural geology offers a significant advantage over traditional deterministic approaches through their capacity to define multiple plausible geological realizations, which is particularly valuable for the ill-posedness of geological reconstruction from often sparse and incomplete data. The use of synthetic data to train a generative model is an opportunity to specify prior knowledge that is vastly more complex than what would be permitted through classical techniques such as regularization. The unconditional models generated have a high degree of geological diversity, capturing variations in topography, layer geometry, and structural features such as dykes and intrusions. This type of synthetic modeling has numerous applications, including data augmentation for training machine learning models, benchmarking geophysical inversion algorithms, and enhancing geostatistical modeling techniques. Our results further reinforce previous studies that highlight the importance of generative approaches in geosciences [30, 41, 35, 36].
Conditional Modeling, Reconstruction, and Uncertainty Quantification.
@@
The ability to generate samples conditioned on sparse borehole data presents a major breakthrough for geological interpolation tasks. The technique can succeed where tools such as kriging often struggle, for example, when borehole data is limited or categorical. The generative model approach produces an ensemble of realizations that not only fit the observed data but also adhere to a probability distribution of plausible geology encoded into the synthetic training set. The generated models are not only consistent with the conditional data but also diverse, assisting tasks such as risk assessment and decision-making in resource exploration, subsurface engineering, and geological hazard analysis. By providing a suite of possible geological scenarios, our approach allows decision-makers to incorporate uncertainty into their models rather than relying on a single deterministic solution.
Implications and Future Directions.
@@
The ability to generate diverse, geologically plausible models has far-reaching implications for a variety of geoscientific applications. One promising direction is the integration of our generative framework with geophysical inversion workflows, where subsurface properties inferred from geophysical data can be used to further constrain model realizations. Additionally, combining our approach with Bayesian inference techniques could enable probabilistic geological modeling, providing a more rigorous quantification of uncertainty.
Conclusions.
@@
The incorporation of generative AI into structural geology represents a paradigm shift, enabling more flexible, data-driven, and probabilistic interpretations of the Earth’s subsurface. As these methods continue to evolve, they have the potential to transform geological modeling, supporting improved resource exploration through, e.g., providing novel types of regularization for geophysical inverse problems, as well as geohazard assessment and geotechnical decision-making.

### v3

@@ 4 Discussion
Our results demonstrate the feasibility and potential of using generative artificial intelligence (AI) models, particularly flow matching techniques, to construct realistic and geologically plausible three-dimensional structural models. The work presented here integrates two key components. First, we construct a process for creating a large, synthetic, and geologically consistent dataset that seeks to embed the rich and complex knowledge of structural geology into a mathematical framework suitable for training machine learning models. Second, we train a generative AI model that implicitly learns both unconditional and conditional probability distributions of geological structures. By doing so, it efficiently captures essential characteristics of largescale lithological units in the Earth’s crust and provides a highly complex prior to tackle the challenges common in geoscience applications, including data sparsity, uncertainty quantification, and the integration of diverse data modalities. Although the underlying simulation engine could, in principle, operate at finer scales or incorporate continuous physical properties such as conductivity, permeability, density, our focus here is on modeling the categorical organization of major lithologic bodies conditioned by surface and borehole constraints.
@@
Advancements in Generative Modeling for Structural Geology. Generative AI for structural geology offers a significant advantage over traditional deterministic approaches through their capacity to define multiple plausible geological realizations, which is particularly valuable for the ill-posedness of geological reconstruction from often sparse and incomplete data. The use of synthetic data to train a generative model is an opportunity to specify prior knowledge that is vastly more complex than what would be permitted through classical techniques such as regularization. The unconditional models generated have a high degree of geological diversity, capturing variations in topography, layer geometry, and structural features such as dykes and intrusions. This type of synthetic modeling has numerous applications, including data augmentation for training machine learning models, benchmarking geophysical inversion algorithms, and enhancing geostatistical modeling techniques. Our results further reinforce previous studies that highlight the importance of generative approaches in geosciences [wellmann2015PynoddyExperimentalPlatform, bergen2019MachineLearningDatadriven, woollam2021seisbenchtoolboxmachine].
@@
Conditional Modeling, Reconstruction, and Uncertainty Quantification. The ability to generate samples conditioned on sparse borehole data presents a major breakthrough for geological interpolation tasks. The technique can succeed where tools such as kriging often struggle, for example, when borehole data is limited or categorical. The generative model approach produces an ensemble of realizations that not only fit the observed data but also adhere to a probability distribution of plausible geology encoded into the synthetic training set. The generated models are not only consistent with the conditional data but also diverse, assisting tasks such as risk assessment and decision-making in resource exploration, subsurface engineering, and geological hazard analysis. By providing a suite of possible geological scenarios, our approach allows decision-makers to incorporate uncertainty into their models rather than relying on a single deterministic solution.
@@
Implications and Future Directions. The ability to generate diverse, geologically plausible models has far-reaching implications for a variety of geoscientific applications. One promising direction is the integration of our generative framework with geophysical inversion workflows, where subsurface properties inferred from geophysical data can be used to further constrain model realizations. Additionally, combining our approach with Bayesian inference techniques could enable probabilistic geological modeling, providing a more rigorous quantification of uncertainty. Future work could also extend this framework to finer resolutions or physical property domains where multiscale variability becomes essential.
@@
Conclusions. The incorporation of generative AI into structural geology represents a paradigm shift, enabling more flexible, data-driven, and probabilistic interpretations of the Earth’s subsurface. As these methods continue to evolve, they have the potential to transform geological modeling, supporting improved resource exploration through, e.g., providing novel types of regularization for geophysical inverse problems, as well as geohazard assessment and geotechnical decision-making.
@@
Acknowledgments
For computer time, this research partially used the resources of the Supercomputing Core Laboratory at King Abdullah University of Science and Technology (KAUST).
Open Research
The StructuralGeo simulation code (version 1.0) used to generate the synthetic geological models in this study is preserved at Zenodo [StructuralGeo_v1] and openly developed at GitHub (https://github.com/eldadHaber/StructuralGeo/releases/tag/v1.0), under a permissive open-source license.
The generative-model training and evaluation code is preserved at Zenodo [FlowTrain_v1] and developed at GitHub
(https://github.com/chipnbits/flowtrain_stochastic_interpolation/releases/tag/v1.0.2). This code supports reproducing both conditional and unconditional experiments.
[…]
All software is released under open-source licenses. No registration is required to access the data or software.
Author Contributions Statement
S.G. developed the synthetic dataset generator and was involved in the generative AI model development. V.O. and S.Z. trained the foundation model. G.T. led the training and together with D.K. organized the KAUST-based effort. E.H. conceived of and led the overall project. All authors participated in preparing and reviewing the manuscript.
Competing Interests Statement: The authors declare no competing interests.
Authorship and AI Tools: ChatGPT (GPT-5, OpenAI) was used in this manuscript to assist with spelling, grammar, and improvements to sentence and paragraph structure. The authors reviewed, verified, and approved all suggestions and take full responsibility for the content.